OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to y = 3x+6 and passes through the point (-10, 2.5)

OpenStudy (anonymous):

Okay... so if a line is parallel to another line, then the slope of the two lines would be the same.

OpenStudy (anonymous):

And if you want to find out the equation of the line, plug in x and y values in the equation. X and Y values are given. In your case, it is (-10, 2.5). Basically, the equation would then be y=3x+b.

OpenStudy (anonymous):

the slope intercept form is (y-y1) = m(x-x1) m=3(because the lines are parallel---> same slope plug in the given values of x and y into x1 and y1 and solve for y (so that y is on the left hand side). This would give your answer!

OpenStudy (anonymous):

Do you get it? Let me know if you don't.

OpenStudy (waleed_imtiaz):

@benfraser1012 I am confused with the question... If I have to find the equation using slope intercept form.... then slope intercept form is y=mx+c not (y-y1) = m(x-x1)...... Slope would be 3

OpenStudy (waleed_imtiaz):

And now I dont know where to put the points... x1 and y1........

OpenStudy (anonymous):

okay. Yes, the slope intercept form would be y=mx+c (like you said) but you don't know what "c" is yet, right. You just know that m=3. So to find out C, you use the formula that i mentioned earlier.

OpenStudy (anonymous):

So it would be y-(-10)=3(x-2.5). y+10=3x-7.5 y=3x-17.5

OpenStudy (waleed_imtiaz):

so from this c=-17.5.. right ?

OpenStudy (anonymous):

yes.

OpenStudy (waleed_imtiaz):

SO is this the equation y=3x-17.5

OpenStudy (anonymous):

Yes. That's your answer.

OpenStudy (waleed_imtiaz):

Okay... now I read the question correctly...... We've to write the equation in slope intercept form using the given data..... So it's quite easy.... just use point-slope form and the convert it into slope intercept form... right ?

OpenStudy (anonymous):

You got it!

OpenStudy (waleed_imtiaz):

Yea.. My mistake I read the question carelessly...... Thanku bro..